NAOJ GW Elog Logbook 3.2

This sample absorption has been remeasured as there was some discrepancy with measurement few years old.
This one agrees with the one I did few months ago : absorption is around 18 ppm
(Pin = 7.316W, Pt = 6.29W, R = 0.8342 ppm/cm)
There are however strange patterns on the XZ YZ measurements that were not present before.

We started evacuation of the BS chamber yesterday. I opened the gate valve between BS and NM2 today.

Fitting this data with a cosine gives input polarization equal to output sample polarization for angle about 41.34 deg and the phase about -6.5 deg

Marc, Matteo, Takahashi-san, Yuhang
We opened BS chamber today to check if there is any issue for the magnet.
We found one magnet falls down as shows in the attached figure.
Takahashi-san will help us prepare a replacement magnet. The replacement work will be done in the next next week.

We have done some more check before opening the BS vacuum chamber.
1. Check BS P and Y transfer functions. They are similar with reference as shows in the fig1 and fig2.
2. Check coupling between P and Y. A substantial coupling increase is found as shows in the fig3.
3. We tried to close BS local control loop. But failed.
The results of test 2 and 3 make us really need to open BS chamber to check. The opening work is scheduled to be in this afternoon.

Today I tried to tweak the input pump power and did several long Z scan to find the 2 surfaces of the sample.
Because it did not succeed at first, I reinstalled the bulk reference sample and got R_bulk = 0.8342 cm/W
Finally, by taking 10 averages and mean I could see the surfaces and started the measurement at Z_center = 62.75 mm
Using the value of this long Z scan the absorption is consistent with the previous measurement :
AC ~ 1.5e-4V
DC~4V
Ptrans = 6.290W
Pinc = 7.315W
gives absorption ~ 22 ppm/cm
Measurement is on-going with 0.5mm step size and 30 mm radius.
Do not touch the sample this week !

Katsuki-san, Marc
Today we removed the AZTEC sample in order to prepare the absorption and birefringence measurement of the TAMA T1 sample.
We reinstalled the steering mirror in the injection part to dump the sample reflection (see fig1)
We installed a flipping mirror before the birefringence readout to redirect the pump beam to the high power power meter (see fig2).
We reconnected the absorption photodiode to the lockin amplifier.
We installed the surface reference sample and with z = 42.3mm and zIU = 66 mm we got R_surf = 18.72 /W
We installed the bulk reference sample with z = 40.95 mm and zIU = 65.7 mm and got R_bulk = 0.6679 cm/W.
We installed the TAMA T1 and using DC values could find Y_center = 121.92 mm and X_center = 326.74 mm.
We increased the pump laser current to 6A (beforehand we changed the HWP angle that control the power to 9 deg that corresponds to the min pump power) which gave Pt ~6.1W.
We realigned the DC at the previous z_center (62.4mm) and did long z scan but could not find clear surfaces absorption neither on the AC nor the phase.
This seems to confirm that this sample absorption is quite low ( previously measured at 18 ppm).
Because it is the first time using such high power with all the birefringence optical component we turned off the laser today and will start the absorption measurement tomorrow morning.

Katsuki-san, Marc
Today we checked the relationship between the input polarization angle and the AZTEC sample polarization angle.
We confirmed that thanks to the realignment the maximal power of both photodiode is reached without sample.
Further analysis to follow.
Fitting this data with a cosine gives input polarization equal to output sample polarization for angle about 41.34 deg and the phase about -6.5 deg

Here are the figures.

During the week-end we took birefringence measurements of the AZTEC sample rotated by 180 deg and back to same orientation as the absorption measurement (ie 0 deg).
Due to some mistakes it was not possible to perform too many measurements (once labview is exited it reset the lockin amplfiier to absorption measurement)..
To avoid confusion I report here the measurement with 180 deg rotation.

Recently, it is found that GRMC loop cannot be closed. To check what is the problem, I disconnected the slow scan of MZ and put MZ PZT offset to a value that GRMC transmission is the highest. After that, according to the original setting, I have done following checks
0. GRMC has a good alignment.
1. PDH signal has 316mV pk-pk checked from EPS1.
2. GRMC has loop sign of INV, which is as design.
3. The RF source phase is reloaded. The phase of RF source is 125deg. When it is changed to 35deg, the signal around resonance becomes flat. This indicates the RF signal phase is still a good one.
4. There is a switch which has +/- sign. This doesn't decide the sign of control loop. But when we use this type of servo for CC1/2 controls, we need to flip this switch. I tried to flip this switch, but it doesn't help to close loop.
5. GRMC transmission is checked to have 1.13V peak. This is two times smaller than the value written by Pierre.
6. Loop gain is 3 as usually used.
7. Threshold for peak identification is -0.55V. This is as required.
8. The GR power reaching AOM is measured to be 44mW, whose nominal value is 50mW.
These checks show little issues but they should not prevent the GRMC locking. More investigation is required.

After the earthquake on 7th this month, we confirmed the normal operation of vacuum system, mirror suspension system. Especially, the filter cavity alignment was recovered and mirror's oplev shows usual spectra.
However, we found already that the driving of BS has problem. The problem is when we try to move yaw or pitch, the beam reflected by BS moves diagonally.
To confirm if there is any issue for magnets, I took a photo of BS magnets. Attached photo shows this check. Two red boxes indicate the location of four coil-magnet actuators. An arrow points to the magnet which seems to be broken. This magnet may be knocked during earthquake and get tilted. So the coil cannot drive it properly.
We have done some more check before opening the BS vacuum chamber.
1. Check BS P and Y transfer functions. They are similar with reference as shows in the fig1 and fig2.
2. Check coupling between P and Y. A substantial coupling increase is found as shows in the fig3.
3. We tried to close BS local control loop. But failed.
The results of test 2 and 3 make us really need to open BS chamber to check. The opening work is scheduled to be in this afternoon.
Marc, Matteo, Takahashi-san, Yuhang
We opened BS chamber today to check if there is any issue for the magnet.
We found one magnet falls down as shows in the attached figure.
Takahashi-san will help us prepare a replacement magnet. The replacement work will be done in the next next week.
We started evacuation of the BS chamber yesterday. I opened the gate valve between BS and NM2 today.

Katsuki-san, Marc
In order to mitigate the misalignment induced by the sample, we decided to move the 2 PDs closer to the waist.
For that, we also needed to change the ODs mounts for a smaller one.
We confirmed that with the realignment the maximal power of s or p polarizations with sample is smaller than the maxima without sample.
We started measurement with s polarization at input and the AZTEC sample rotated by 180 deg

Just to clarify :
There is some input polarization angle for which the the s polarization power is higher with sample than the maximum without sample.
This hints for either scattered light but more probably some misalignment (maybe the beam hits the side of the PDs that fakes strong power?)

Shinkosha evaluation plate # 10 is available inside PCI clean room.
Documentation is below the PC.

After spending time trying to use another PC, another usb hub, another usb cable and others Kinesis version that all failed I installed a spare motorized HWP and confirmed that it worked perfectly fine.
It is now possible to control perfectly fine the 2 motorized HWPs at the same time.

Michael and Yuhang
We checked filter cavity alignment yesterday. The pico-motor of PR/BS and END mirrors are moved. The movement of PR is to recover its reference on BS chamber. The movement of BS is to recover its reference in filter cavity transmission camera. The movement of END mirror is to recover the flash.
After these movement, we checked the oplev spectrum as attached in this elog. They look fine. So this time, no touching issue of mirror was found.

Katsuki-san, Marc
This morning we went to PCI to check the ATEC sample.
We found out that the peak visible in the polarization angle measurement is likely due to a dust so we removed it. Except that the sample seems perfectly fine.
We brought a laptop with Kinesis installed and connected it to the motorized HWP.
We got the same issue as with the PC (only jog by 1 deg).
We tried to home it and will check after lunch.
If problems persist we will contact Thorlabs and switch the 2 motorized HWP as the other is working fine.

This entry reports the birefringence measurement of the AZTEC sample with no rotation,
Note that a quite strong earthquake happend right after the measurement with input polarization = 0 deg...
Note that in that case the input polarization angle of 0 deg is made with hwp angle = 346 deg..
While the mean value of the polarization angle measurement seems reasonable, there is now a strange 'peak' on the top right of the sample...
We'll go to PCI to check if the sample is fine.

In order to further investigate the effects of the input beam incident angle we mesured the polarization angle as a function of the input beam polarization angle.
Data are attached to this entry.
Note that :
There are input polarization angle for which the s polarization power transmitted with sample is higher than transmitted s polarization power without sample..
The minimum s polarization seems to be reached for hwp angle = 71 deg (ie 26 +45 deg) and using this polarization to normalize our signals means that the hwp angle = 341 deg that was assumed to be fully s polarization is in fact 10 deg polarization angle...
Just to clarify :
There is some input polarization angle for which the the s polarization power is higher with sample than the maximum without sample.
This hints for either scattered light but more probably some misalignment (maybe the beam hits the side of the PDs that fakes strong power?)